Consider the following information about a simple country that produces two different goods:

Year

Price Milk

Quantity Milk

Price Cereal

Quantity Cereal

2014

$4

200

$5

300

2015

$4.25

250

$5.50

330

2016

$4.50

300

$6.00

350

From 2014 to 2015 the growth rate of nominal GDP was ____% and the growth rate of real GDP was _____%.

Enter a number rounded to two decimal places. Assume 2014 was the base year.
Answer #1

NGDP in 2014 = 4*200+5*300 = 2300

NGDP in 2015 = 4.25*250+5.5*330 = 2877.5

Growth rate of NGDP = (2877.5-2300)/2300*100 = 25.11%

RGDP in 2014 = 2300

RGDP in 2015 = 4*250+5*330 = 2650

Growth rate of RGDP = (2650-2300)/2300*100 = 15.22%